OpenStudy (anonymous):

WILL FAN AND MEDAL The sum of two numbers is 22, and their difference is 12. What are the numbers?

OpenStudy (anonymous):

10+12=22 22-12=10 22-10=12

OpenStudy (emma.davenport):

18+4=22 18-4=12